# A Step Toward World Models: A Survey on Robotic Manipulation

### v1

@@ Abstract
Autonomous agents are increasingly expected to operate in complex, dynamic, and uncertain environments, performing tasks such as manipulation, navigation, and decision-making. Achieving these capabilities requires agents to understand the underlying mechanisms and dynamics of the world, moving beyond purely reactive control or simple replication of observed states. This motivates the development of world models as internal representations that encode environmental states, capture dynamics, and enable prediction, planning, and reasoning. Despite growing interest, the definition, scope, architectures, and essential capabilities of world models remain ambiguous. In this survey, rather than directly imposing a fixed definition and limiting our scope to methods explicitly labeled as world models, we examine approaches that exhibit the core capabilities of world models through a review of methods in robotic manipulation. We analyze their roles across perception, prediction, and control, identify key challenges and solutions, and distill the core components, capabilities, and functions that a real world model should possess. Building on this analysis, we aim to outline a roadmap for developing generalizable and practical world models for robotics. This is an initial version of the survey. The content will be expanded and refined in future updates.
@@ I Introduction
Robots are designed to extend human capabilities, assisting in tasks that are dangerous, repetitive, or demand high precision, thereby enhancing productivity across diverse applications. Achieving such capabilities requires agents to move beyond reactive control and the mere replication of observed states, instead developing the ability to model, reason about, and predict environmental dynamics. In this context, world models have emerged as powerful internal representations that enable robots to anticipate future outcomes, support effective decision-making, and ultimately act intelligently in the real world. Richens et al. [1] argue that any agent capable of generalizing to solve multi-step tasks must implicitly learn a predictive model of its environment, e.g., a world model.
@@
The concept of “world models” in computer science dates back to the 1960s [2], and numerous methods have since been proposed as steps toward more capable models [3, 4, 5, 6, 7, 8, 9, 10], although not all of these works explicitly identify themselves as world models. For example, Wang et al. [9, 10, 5, 6, 7, 8] leverage video generation models as a form of world models, which encode extensive world knowledge from large-scale training data and can predict future states based current observations and/or actions. LeCun et al. [11, 12, 13, 14, 15] emphasizes modeling abstract world state representations, while Zitkovich et al. [16, 17, 18] utilize vision-language-action models (VLA) models that do not explicitly generate future states. The scope of existing methods varies from 2D scene prediction to 4D world modeling [19, 20, 21, 22, 23, 24], reflecting different understandings of what it means to model the world. The observation viewpoint of the world includes both third-person (exocentric) [21, 25, 26] and first-person (egocentric) [27, 28] perspectives.
@@
World models play a critical role in robotic learning in two ways. They allow robots to improve autonomous policies by simulating multiple action proposals and selecting the optimal one for execution [29, 12, 13, 14, 30, 21]. They also support scalable policy training and evaluation by generating realistic rollouts and physical interactions, providing an efficient alternative to collecting data in the real world. [16, 31, 32, 33, 34, 35]. From a functional standpoint, current approaches range from single-purpose models, such as those designed for visual planning [30, 36], future-scene generation [37, 33, 38], or action prediction [39], to more integrated systems that couple multiple abilities within a unified framework [40, 27, 41, 42, 43].
@@
  • Is human cognition [44, 45] the ultimate goal for world models?
@@
  • Comprehensive taxonomy of world model architectures. We provide a systematic categorization of existing designs, including latent space modeling methods, video generation-based models, direct projection based methods and other emerging structures.
  • Functional analysis. We discuss the diverse roles of world models in robotics, including robotic learning, evaluation, and planning, highlighting their contribution to autonomous control.
@@
  • Challenges and future directions. We summarize key challenges, including scalability, physical awareness, and generalization, and discuss potential research directions and solutions toward building practical, real-world models.
Related Surveys. Our survey differs substantially from existing reviews. Several surveys have examined world models in robotics, but most focus on specific aspects and provide limited conceptual analysis. For example, Yu et al. [46] emphasize video generation, Kong et al. [47] cover 3D/4D world modeling, Ai et al. [48] study dynamics learning, and Lin et al. [49] address physics cognition. Long et al. [50] review architectures and functional roles of world models, whereas Zhu et al. [51, 52, 53] primarily compile representative works. While these surveys provide valuable overviews, they offer limited discussion of the essential characteristics and functional requirements of comprehensive world models for embodied agents. In contrast, our survey presents a holistic, problem-centered perspective, highlighting key challenges, solution strategies, and future directions for world modeling in robotics.
@@
  • Prediction Tasks: AP: Action Prediction, PL: Policy Learning, VP: Visual Planning, Static: Static Visual Prediction, underlineAction-cond.: Action-conditioned Visual Prediction.
  • Input & Output: L: Language, V: Video, A: Action, S: State, I: Image, P: Point Cloud, Tr: Trajectory, Ar: Autoregressive
  • Core Components: CLIP: Contrastive Language-Image Pre-training, DiT: Diffusion Transformer, IDM: Inverse Dynamics Model, GPT: Generative Pre-trained Transformer, LLM: Large Language Model, LSTM: Long Short-Term Memory, RSSM: Recurrent State-Space Model, U-Net: U-shaped Convolutional Neural Network, VAE: Variational AutoEncoder, VDM: Video Diffusion Model, ViT: Vision Transformer, VLM: Vision-Language Model, VQ: Vector Quantization.
@@
PlaNet [29] V+A V+A A RSSM
DreamerV1,V2,V3 [12, 13, 14, 15] V+A V+A A RSSM
PaLM-E [54] V+L V+L A LLMs, ViTs
OpenVLA [55] V+L+A V+L A VLA
Plan2Explore [39] V+A V+A A RSSM
FOCUS [25] V+A V+A A RSSM
EgoAgent [27] V V+A A JEPA
GR00T  [44] V+L+A+S V+L+A+S A VLA, VLM
THICK [56] V+A V+A A Hierarchical, RSSM
DayDreamer [14] V+A V+A A RSSM
RetryingVisualMPC [30] V I A Registration network
Genie [20] V I/V V VQ-VAE, VQ-VAE
GE [57] V + L L+I V, Ar VAE, DiT, CLIP
GR-2 [58] V V+L+S V+A+S, Ar CLIP, VQGAN, GPT, condition VAE
UniSim [19] V+L+A V+A V Stable Diffusion
UniPi [59] V+L L+I V IDM, VDM
DiVA [9] V+L V+L V+A VDM, CLIP, VAE
WorldVLA [40] V+L V+L+A V+A, Ar VLA, VQ-GAN
iVideoGPT [6] V+A V+A V, Ar VQ-GAN, GPT
DreamVLA [60] V+L+A+S V+L+S A, Ar CLIP, Masked Autoencoders, LLMs (GPT-2), DiT
Seer [61] V+L+A+S V+L+S V+A, Ar CLIP, ViT, GPT
EnerVerse [8] V+L+A V+L V, Ar VAE, VDM
SayCan [62] V+L+A V+L A LLMs
FOCUS [25] I+A+S I+A+S I+A RSSM
Visual MPC [63] V+A+S V+A+S V+A LSTM
CDNA [36] V+A+S V+A+S V LSTM
VisualForesight [64] I V+goal(I/L) A LSTM
VLP [5] I+V+L+A I+L V+A VLM, VDM
FLIP [65] V+L V+L V CVAE, DiT, LIV
RoboDreamer [66] V+L V+L V Text Parser, VDM, IDM
COMBO [7] V+L+A V+A V+A VDM, VLM, Tree Search
AVID [67] V+A V+A V VDM, Action Adapter
SuSIE [68] V+L+A V+L V+A Image-Editing Diffusion Model
3D-VLA [41] V+L+A+S V+L V+A+P 3D-LLM, Stable Diffusion
AETHER [23] V V+A V Geometric Encoding, VDM
TesserAct [24] V+L+A V+L V+A VDM
IRASim [43] V+A V+Tr V, Ar DiT
3DFlowAction [69] V+L V+L+P Tr, Ar VDM, CLIP
THICK [56] V+A+S V+A V+A+S Context-Specific RSSM
FlowDreamer [21] V+A V+A V U-Net, Stable Diffusion
DREMA [33] V+A+S V+S V+A Gaussian Splatting
AVDC [70] V+L V+L V+Optical Flow+A U-Net, CLIP-Text, Optical Flow Model
GWM [34] V+A V+A V 3D Gaussian VAE, DiT
Plan2Explore [39] V V A DM, Dreamer
DreamGen [38] V V+L V DiT
UWM [71] V V+A V+P IDM, DiT
CLOVER[22] V+L V+L V+A ViTs, CLIP, VDM
Pandora [72] V+L V+L V, Ar LLM, DynamiCrafter Model
PAR [42] V+L V+L+A V+A, Ar 3D-VAE, Phi Model, DiT, Causal Transformer
RoboEnvision [10] V+L V+L V+A VLM, DiT
HMA [73] V+L V+L V+A VLM, DiT
WorldGym [74] V+A V+A V VLM, DiT
VideoWorld [75] V V V+A, Ar IDM, VQ-VAE, Ar Transformer
PlaySlot [26] V V V, Ar IDM, Recursive Encoder-Decoder, Ar Transformer
LOReL [76] V+L V+A V+A LSTM
V-JEPA 2 [77] V+A V+A V+A JEPA, LLM
UP-VLA [78] V+L V+L V+A+L VLA, Decoder
CoT-VLA [79] V+L V+L V+A VLA, Decoder
@@ II-A What Is the “World” to Be Modeled?
Despite the debate among philosophers about the ultimate nature of the world, the world can be roughly described as a set of entities, each with its own attributes or properties, along with the relationships and interactions that connect them. These attributes, such as shape, size, material, or state, and the connections, which can be spatial, causal, functional, or temporal, mean that objects, agents, and features are not only statically arranged but also evolve and influence one another over time. In order to interact effectively with such a world, an intelligent agent must capture critical information about entities, their properties, and their interactions. Collectively, these entities and interactions form a rich and dynamic environment in which an agent must actively explore, interact, and learn to achieve its goals. This naturally raises the question of what fundamental capabilities underpin an agent’s ability to capture and reason about such complex dynamics, as well as what forms of representation, learning, and interaction are required to model and act within an uncertain and evolving world.
@@ II-B World Models Empowering Robot Intelligence
Embodied intelligence refers to a system’s ability to perceive, reason, and learn through direct interaction with its environment. Unlike traditional AI confined to abstract or symbolic domains, embodied intelligence integrates a physical body, sensors, actuators, and computational processes that together enable situated perception, reasoning, and action. Intelligent robotic agents serve as the primary physical instantiation of embodied intelligence. They inherently combine perception (via sensors), cognition (for learning and reasoning), and motor control (via actuators) to operate autonomously and acquire knowledge from real-world experience, much like biological organisms.
However, because intelligent agents perceive only a partial and noisy projection of reality through their sensors, many underlying relationships and causal dependencies remain latent. This limitation makes structured internal representations essential for prediction, planning, and multi-step reasoning. To achieve robust and efficient embodied intelligence, recent research introduces the notion of world models, which serve as internal representations that capture environmental dynamics and common-sense regularities of how the world operates. By internally simulating potential outcomes, world models empower embodied agents to understand their context, anticipate the consequences of actions, and plan complex behaviors before execution, thereby reducing reliance on costly or unsafe real-world trial and error.
@@ II-C Competing Perspectives on World Models
Although the concept of a “world model” is prevalent in computer science, its definition remains unsettled, with ongoing debate in the research community regarding its fundamental nature and role in intelligence [80]. A central point of contention concerns its generative capability, as illustrated by NVIDIA [81], who define world models as systems that learn environmental dynamics from multimodal data and generate videos capturing spatial and physical properties. Emphasizing action dependence, Sudhakar et al. [37, 40] characterize world models specifically as action-conditioned video generation models, distinguishing them from conventional video prediction. Similarly, Hafner et al. [14, 12, 13, 15, 82] identify action-conditioned prediction as a core feature of world models, emphasizing the prediction of latent representations rather than raw observations. Despite these differing perspectives, a common consensus emerges: world models aim to construct internal representations that capture environmental dynamics and action consequences, thereby enabling the prediction of future states.
Figure 1: A visualization of an agent [11], where the world model predicts possible future world states as a function of imagined actions sequences proposed by the actor.

### v2

@@ Abstract
Autonomous agents are increasingly expected to operate in complex, dynamic, and uncertain environments, performing tasks such as manipulation, navigation, and decision-making. Achieving these capabilities requires agents to understand the underlying mechanisms and dynamics of the world, moving beyond reactive control or simple replication of observed states. This motivates the development of world models as internal representations that encode environmental states, capture dynamics, and support prediction, planning, and reasoning. Despite growing interest, the definition, scope, architectures, and essential capabilities of world models remain ambiguous. In this survey, we go beyond prescribing a fixed definition and limiting our scope to methods explicitly labeled as world models. Instead, we examine approaches that exhibit the core capabilities of world models through a review of methods in robotic manipulation. We analyze their roles across perception, prediction, and control, identify key challenges and solutions, and distill the core components, capabilities, and functions that a fully realized world model should possess. Building on this analysis, we aim to motivate further development toward generalizable and practical world models for robotics. This is an initial version of the survey. The content will be expanded and refined in future updates.
@@ I Introduction
Figure 1: Conceptual flow of the survey. The survey aims to clarify the motivation behind world modeling, explore its essential scope and development, and illuminate pathways toward more general and capable world models.
@@
Autonomous agents are designed to extend human capabilities, assisting in tasks that are dangerous, repetitive, or demand high precision, thereby enhancing productivity across diverse applications. Robots serve as their physically embodied form, aiming to realize seamless interaction with the physical world. Achieving such capabilities requires agents to move beyond reactive control and the mere imitation of observed behaviors, instead developing the ability to model, reason about, and predict environmental dynamics. In this context, world models have emerged as powerful internal representations that enable robots to anticipate future outcomes, plan effectively, and ultimately act intelligently in the real world. Richens et al. [richens2025general] argue that any agent capable of generalizing to solve multi-step tasks must implicitly learn a predictive model of its environment, e.g., a world model.
@@
The concept of “world models” in computer science dates back to the 1960s [minsky1975framework], and numerous methods have since been proposed toward more capable models [schmidhuber2015learning, ha2018world, du2023video, wu2024ivideogpt, zhang2025combo, huang2025enerverse, wang2025language, yang2025roboenvision], though not all explicitly identify as world models. For example, Wang et al. [wang2025language, yang2025roboenvision, du2023video, wu2024ivideogpt, zhang2025combo, huang2025enerverse] leverage video generation models as a form of world models, which encode extensive world knowledge from large-scale training data and can predict future states based current observations and/or actions. LeCun et al. [lecun2022path, hafnerdream, hafner2021mastering, wu2023daydreamer, hafner2023mastering] emphasizes modeling abstract world state representations, while Zitkovich et al. [zitkovich2023rt, huang2024embodied, hong2024multiply] utilize vision-language-action models (VLA) models that do not explicitly generate future states. The scope of existing methods varies from 2D scene prediction to 4D world modeling [yang2023learning, bruce2024genie, guo2025flowdreamer, bu2024closed, team2025aether, zhen2025tesseract], reflecting different understandings of what it means to model the world. The observation viewpoint of the world includes both third-person (exocentric) [guo2025flowdreamer, ferraro2025focus, villar2025playslot] and first-person (egocentric) [chen2025egoagent, grauman2024ego] perspectives.
@@
World models play a critical role in robotic learning in two ways. They allow robots to improve autonomous policies by the internal prediction and simulation of environmental dynamics, future states, and/or action outcomes [hafner2019learning, hafnerdream, hafner2021mastering, wu2023daydreamer, ebert2018robustness, guo2025flowdreamer]. They also support scalable policy training and evaluation by generating realistic rollouts and physical interactions [zitkovich2023rt, black2024pi_0, team2025gemini, barcellona2025dream, lu2025gwm, liu2025rdt]. From a functional standpoint, current approaches range from single-purpose models, such as those for visual planning [ebert2018robustness, finn2016unsupervised], future-scene generation [sudhakar2024controlling, barcellona2025dream, jang2025dreamgen], or action prediction [sekar2020planning], to more integrated systems that couple multiple abilities within a unified framework [cen2025worldvla, chen2025egoagent, zhen20243d, song2025physical, zhu2025irasim].
@@
  • Is human cognition [bjorck2025gr00t, wang2025dmwm] the ultimate goal for world models?
@@
  • Comprehensive taxonomy of world model architectures. We provide a systematic categorization of existing designs, including latent space modeling methods, video generation-based models, direct projection based (implicit modeling) methods and other emerging structures.
  • Functional analysis. We discuss diverse roles of world models in robotic manipulation, including action prediction and planning, robotic learning, and evaluation, highlighting their contribution to intelligent manipulation.
@@
  • Challenges and future directions. We summarize key challenges, including long-horizon reasoning, physical awareness, and generalization, and discuss potential research directions and solutions toward building practical, real-world models.
Related Surveys. Our survey differs substantially from existing reviews. Several prior surveys have examined world models in robotics, but most focus on specific aspects and provide limited conceptual analysis. For example, Yu et al. [yu2025survey] emphasize video generation, Kong et al. [kong20253d] cover 3D/4D world modeling, Ai et al. [ai2025review] study dynamics learning, and Lin et al. [lin2025exploring] address physics cognition. Long et al. [long2025survey] review architectures and functional roles of world models, whereas Zhu et al. [zhu2024sora, liang2025large, ding2025understanding] primarily compile representative works. While these surveys provide valuable overviews, they offer limited discussion of the essential characteristics and functional requirements of comprehensive world models. In contrast, our survey presents a holistic, problem-centered perspective, highlighting key challenges, solution strategies, and future directions for world modeling in robotics.
@@
  • Prediction Tasks: AP: Action Prediction, PL: Policy Learning, VP: Visual Planning, Static: Static Visual Prediction, Action-cond.: Action-conditioned Visual Prediction.
  • Input & Output: L: Language, V: Video, A: Action, S: State, I: Image, P: Point Cloud, Tr: Trajectory.
  • Core Components: CLIP: Contrastive Language-Image Pre-training, DiT: Diffusion Transformer, IDM: Inverse Dynamics Model, GPT: Generative Pre-trained Transformer, LLM: Large Language Model, LSTM: Long Short-Term Memory, RSSM: Recurrent State-Space Model, U-Net: U-shaped Convolutional Neural Network, VAE: Variational AutoEncoder, VDM: Video Diffusion Model, ViT: Vision Transformer, VLM: Vision-Language Model, VQ: Vector Quantization, Ar: Autoregressive.
@@
PlaNet [hafner2019learning] V+A V+A A RSSM
DreamerV1,V2,V3 [hafnerdream, hafner2021mastering, wu2023daydreamer, hafner2023mastering] V+A V+A A RSSM
PaLM-E [driess2023palm] V+L V+L A LLMs, ViTs
OpenVLA [kim2025openvla] V+L+A V+L A VLA
Plan2Explore [sekar2020planning] V+A V+A A RSSM
FOCUS [ferraro2025focus] V+A V+A A RSSM
EgoAgent [chen2025egoagent] V V+A A JEPA
GR00T  [bjorck2025gr00t] V+L+A+S V+L+A+S A VLA, VLM
THICK [gumbsch2023learning] V+A V+A A Hierarchical, RSSM
DayDreamer [wu2023daydreamer] V+A V+A A RSSM
RetryingVisualMPC [ebert2018robustness] V I A Registration network
Genie [bruce2024genie] V I/V V VQ-VAE, VQ-VAE
GE [liao2025genie] V + L L+I V, Ar VAE, DiT, CLIP
GR-2 [cheang2024gr] V V+L+S V+A+S, Ar CLIP, VQGAN, GPT, condition VAE
UniSim [yang2023learning] V+L+A V+A V Stable Diffusion
UniPi [du2023learning] V+L L+I V IDM, VDM
DiVA [wang2025language] V+L V+L V+A VDM, CLIP, VAE
WorldVLA [cen2025worldvla] V+L V+L+A V+A, Ar VLA, VQ-GAN
iVideoGPT [wu2024ivideogpt] V+A V+A V, Ar VQ-GAN, GPT
DreamVLA [zhang2025dreamvla] V+L+A+S V+L+S A, Ar CLIP, Masked Autoencoders, LLMs (GPT-2), DiT
Seer [tian2025predictive] V+L+A+S V+L+S V+A, Ar CLIP, ViT, GPT
EnerVerse [huang2025enerverse] V+L+A V+L V, Ar VAE, VDM
SayCan [ahn2022can] V+L+A V+L A LLMs
FOCUS [ferraro2025focus] I+A+S I+A+S I+A RSSM
Visual MPC [finn2017deep] V+A+S V+A+S V+A LSTM
CDNA [finn2016unsupervised] V+A+S V+A+S V LSTM
VisualForesight [ebert2018visual] I V+goal(I/L) A LSTM
VLP [du2023video] I+V+L+A I+L V+A VLM, VDM
FLIP [gao2024flip] V+L V+L V CVAE, DiT, LIV
RoboDreamer [zhou2024robodreamer] V+L V+L V Text Parser, VDM, IDM
COMBO [zhang2025combo] V+L+A V+A V+A VDM, VLM, Tree Search
AVID [rigter2025avid] V+A V+A V VDM, Action Adapter
SuSIE [black2024zero] V+L+A V+L V+A Image-Editing Diffusion Model
3D-VLA [zhen20243d] V+L+A+S V+L V+A+P 3D-LLM, Stable Diffusion
AETHER [team2025aether] V V+A V Geometric Encoding, VDM
TesserAct [zhen2025tesseract] V+L+A V+L V+A VDM
IRASim [zhu2025irasim] V+A V+Tr V, Ar DiT
3DFlowAction [zhi20253dflowaction] V+L V+L+P Tr, Ar VDM, CLIP
THICK [gumbsch2023learning] V+A+S V+A V+A+S Context-Specific RSSM
FlowDreamer [guo2025flowdreamer] V+A V+A V U-Net, Stable Diffusion
DREMA [barcellona2025dream] V+A+S V+S V+A Gaussian Splatting
AVDC [ko2024learning] V+L V+L V+Optical Flow+A U-Net, CLIP-Text, Optical Flow Model
GWM [lu2025gwm] V+A V+A V 3D Gaussian VAE, DiT
Plan2Explore [sekar2020planning] V V A DM, Dreamer
DreamGen [jang2025dreamgen] V V+L V DiT
UWM [zhu2025unified] V V+A V+P IDM, DiT
CLOVER[bu2024closed] V+L V+L V+A ViTs, CLIP, VDM
Pandora [xiang2024pandora] V+L V+L V, Ar LLM, DynamiCrafter Model
PAR [song2025physical] V+L V+L+A V+A, Ar 3D-VAE, Phi Model, DiT, Causal Transformer
RoboEnvision [yang2025roboenvision] V+L V+L V+A VLM, DiT
HMA [wang2025learning] V+L V+L V+A VLM, DiT
WorldGym [quevedo2025evaluating] V+A V+A V VLM, DiT
VideoWorld [ren2025videoworld] V V V+A, Ar IDM, VQ-VAE, Ar Transformer
PlaySlot [villar2025playslot] V V V, Ar IDM, Recursive Encoder-Decoder, Ar Transformer
LOReL [nair2022learning] V+L V+A V+A LSTM
V-JEPA 2 [assran2025v] V+A V+A V+A JEPA, LLM
UP-VLA [zhang2025up] V+L V+L V+A+L VLA, Decoder
CoT-VLA [zhao2025cot] V+L V+L V+A VLA, Decoder
@@ II-A What Is the “World” to Be Modeled?
Despite the debate among philosophers about the ultimate nature of the world, the world can be roughly described as a set of entities, each with its own attributes or properties, along with the relationships and interactions that connect them. These attributes, such as shape, size, material, or state, and the connections, can be spatial, causal, functional, or temporal. As a result, objects, agents, and features are not only statically arranged but also evolve and influence one another over time. In order to interact effectively with such a world, an intelligent agent must capture critical information about entities, their properties, and their interactions across spatial and temporal dimensions. Collectively, these entities and interactions form a rich and dynamic environment in which an agent must actively explore, interact, and learn to achieve its goals. This naturally raises the question of what fundamental capabilities underpin an agent’s ability to capture and reason about such complex dynamics, as well as what forms of representation, learning, and interaction are required to model and act within an uncertain and evolving world.
@@ II-B World Models Empowering Robot Intelligence
Embodied intelligence refers to a system’s ability to perceive, reason, and learn through direct interaction with its environment. Unlike traditional AI confined to abstract or symbolic domains, embodied intelligence integrates a physical body, sensors, actuators, and computational processes that jointly enable situated perception, reasoning, and action. Intelligent robotic agents serve as the primary physical instantiation of embodied intelligence. They inherently combine perception (via sensors), cognition (for learning and reasoning), and motor control (via actuators) to operate autonomously and acquire knowledge from real-world experience, much like biological organisms.
However, because intelligent agents perceive only a partial and noisy projection of reality through their sensors, many underlying relationships and causal dependencies remain latent. This limitation makes structured internal representations essential for prediction, planning, and multi-step reasoning. To achieve robust and efficient embodied intelligence, recent research introduces the notion of the “world model”, which serve as internal representations that capture environmental dynamics and common-sense regularities of how the world operates. By internally simulating potential outcomes, world models empower embodied agents to understand their context, anticipate the consequences of actions, and plan complex behaviors before execution, thereby reducing reliance on costly or unsafe real-world trial and error.
@@ II-C Competing Perspectives on World Models
Although the concept of the “world model” is prevalent in computer science, its definition remains unsettled, with ongoing debate in the research community regarding its fundamental nature and role in intelligence [xing2025critiques]. A central point of contention concerns its generative capability, as illustrated by NVIDIA [nvidia_wm2025], who define world models as systems that learn environmental dynamics from multimodal data and generate videos capturing spatial and physical properties. Emphasizing action dependence, Sudhakar et al. [sudhakar2024controlling, cen2025worldvla] characterize world models specifically as action-conditioned video generation models, distinguishing them from conventional video prediction. Similarly, Hafner et al. [wu2023daydreamer, hafnerdream, hafner2021mastering, hafner2023mastering, chen2025vlwm] identify action-conditioned prediction as a core feature of world models, emphasizing the prediction of latent representations rather than raw observations. Despite these differing perspectives, a common consensus emerges: world models aim to construct internal representations that capture environmental dynamics and action consequences, thereby enabling the prediction of future states.
Figure 2: A visualization of an agent [lecun2022path], where the world model predicts possible future world states as a function of imagined actions sequences proposed by the actor.